# Surprisingly High Redundancy in Electronic Structure Data

### v1

@@ Abstract
Machine Learning (ML) models for electronic structure rely on large datasets generated through expensive Kohn-Sham Density Functional Theory simulations. This study reveals a surprisingly high level of redundancy in such datasets across various material systems, including molecules, simple metals, and complex alloys. Our findings challenge the prevailing assumption that large, exhaustive datasets are necessary for accurate ML predictions of electronic structure. We demonstrate that even random pruning can substantially reduce dataset size with minimal loss in predictive accuracy, while a state-of-the-art coverage-based pruning strategy retains chemical accuracy and model generalizability using up to $100$-fold less data and reducing training time by threefold or more. By contrast, widely used importance-based pruning methods, which eliminate seemingly redundant data, can catastrophically fail at higher pruning factors, possibly due to the significant reduction in data coverage. This heretofore unexplored high degree of redundancy in electronic structure data holds the potential to identify a minimal, essential dataset representative of each material class.
@@ I Introduction
The calculation and analysis of the electronic structure of materials underpins much of modern computational materials science. From predicting thermodynamic stability and charge transport to modeling chemical reactivity, the electronic structure serves as a fundamental description of how interacting atoms and electrons give rise to observable properties. Consequently, electronic structure calculations often serve as an important building block of multiscale materials modeling [1] with a significant share of contemporary high performance computing resources being devoted to such simulations [2, 3, 4].
@@
Kohn-Sham Density Functional Theory (KS-DFT) serves as the primary workhorse of modern electronic structure calculations [5, 6, 7]. Although formulated in terms of the so-called Kohn-Sham orbitals, the fundamental unknown in KS-DFT is the electron density. Indeed, many ground state material properties, e.g. structural parameters and elastic constants, may be computed from the electron density field, and it also serves as the starting point for ab initio descriptions of excited state phenomena. However, practical calculations using KS-DFT, especially for complex systems such as disordered alloys or materials with extended defects, can be limited by the cubic scaling computational cost of the method with respect to system size. As such, a considerable share of research has been devoted to the development of specialized solution techniques and high performance computing algorithms to overcome this bottleneck [8, 9, 10, 11, 12, 13, 14, 15, 16].
@@
In recent years, Machine Learning (ML) has enabled the development of powerful surrogate models that can enable the prediction of the ground state electron density and related properties at a fraction of the computational cost of conventional first principles simulations [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. These ML models are typically trained on KS-DFT data and use descriptors of the atomic environment to produce the electron density as the output. Recent studies have demonstrated the success of such ML models in predicting electron densities across a wide range of systems, including pure metals, alloys, insulators, organic molecules and nanostructures  [19, 29, 20, 27, 28, 26]. Moreover, these models show promising extrapolation capabilities, accurately predicting electron densities for unseen compositions in complex alloys, atomic configurations with defects and dislocations, and for chemical species beyond training [29, 28, 27]. The ML models have also been continuously improved to address several challenges, including the enforcement of physical symmetries, incorporation of uncertainty quantification capabilities, and improvements in training and prediction efficiency [18, 30, 27].
@@
The paradigm shift brought on by the development of ML based electronic structure techniques has motivated the creation of extensive datasets [31, 32, 33, 34] — sometimes comprising millions of structures — under the implicit assumption that large, exhaustive sampling is essential to achieve high predictive accuracy and generalizability [35, 36, 37]. While this assumption has driven progress in materials informatics [38, 39, 40], it also raises questions about the efficiency and necessity of such data collection efforts, particularly given the computational cost and environmental footprint associated with high-throughput quantum mechanical calculations.
@@
In this work, we examine a fundamental but largely overlooked question: How much of this electronic structure data is genuinely essential? To answer this, we systematically probe the redundancy in electron density data spanning molecular systems, elemental metals, and complex alloys, and show that a surprisingly small subset is often sufficient to train ML models to within chemical accuracy. Our findings challenge the prevailing view that predictive fidelity and generalizability requires vast amounts of data and suggest that it is possible to identify minimal datasets that dramatically reduce computational cost without sacrificing accuracy. Thus our exploration of the redundancy of electronic structure data can potentially lead to a small foundational dataset for each material class, which in turn can be be used for benchmarking and training foundation models.
@@
The exploration of redundancy in chemical datasets has received prior attention in the literature [41, 42, 43]. For instance Li et. al. [41] have demonstrated a high degree of redundancy in datasets of scalar material properties such as the formation energy and band gaps, and have proposed an iterative strategy to reduce the size of training data. However, there has not been a systematic exploration of redundancies in the electronic structure itself, particularly in the context of ML models. Notable prior efforts include the observation that electronic structure data for various simple systems often lie on low-dimensional manifolds, enabling reconstruction via low-rank tensors [44, 45] or a small number of principal component modes [30, 26]. Somewhat relatedly, recent contributions have also exploited redundancies in the electron density to develop strategic sampling schemes for grid-based representations [46, 47], although the effectiveness of such approaches may depend sensitively on the manner in which core states are modeled. Thus, although previous studies have addressed redundancy in various contexts, statistically grounded one-shot pruning methods — free from ad hoc heuristics — have yet to be explored for electronic structure data. Accordingly, determining a statistically justifiable level of data redundancy remains an open problem, which we address. Thus, the techniques explored here are applicable to any class of material without the need for domain knowledge and possibly even to fields beyond the electron density.
@@
The observation of redundancy in electronic structure data is closely connected to Kohn’s nearsightedness principle [48, 49]. Indeed, this result has been used to deduce that the single particle density matrix that appears in electronic structure calculations can be very sparse [50, 51], and has led to a large number of electronic structure codes that exploit it [8, 52, 53, 54, 55, 56, 57, 58, 9, 59]. In the realm of ML-based electronic structure predictions, similar considerations can be used to motivate the prunability of electronic structure data: since local environments largely determine the electronic structure, many configurations differing only in distant parts of the system are expected to have nearly identical electronic descriptors. This implies substantial redundancy in datasets that exhaustively enumerate configurations, especially in large and complex systems. However, there has not been a systematic exploration of this aspect, particularly in the context of state-of-the-art neural network based ML models of electronic structure.
@@
Neural network performance generally improves with data and model size, often following a power law scaling [60, 61]. However, beyond a certain level, marginal accuracy gains require significantly more data [62, 63]. Thus, power law scaling suggests that much of the training data is redundant and can be pruned without notable performance loss [62].Commonly used approaches for selecting a smaller, representative subset of a large dataset that preserves ML model performance are referred to as data pruning techniques or coreset selection methods. These techniques are often broadly categorized into the following three major approaches [64], importance score based methods [65, 66, 64, 67], coverage driven methods [68, 69, 70] and optimization-based methods [71, 72, 73, 74, 75].
@@
Most existing pruning methods have been evaluated primarily on classification problems, whereas electronic structure prediction is a regression task. Therefore, identifying the degree of prunability of electronic structure data, as well as the most suitable pruning technique remain unexplored. In the present work, two state-of-the-art methods: GraNd [65] and Coverage-Centric Coreset Selection (CCS) [68] are adopted. Optimization-based methods are not considered in this work because, despite their strong theoretical foundation and impressive performance, they are computationally quite expensive, and can be challenging to implement on large-scale datasets [76]. The GraNd technique is adopted because of its intuitive approach, ease of implementation, and ability to identify important examples early in the training. GraNd, an importance-score based method, assigns lower scores to easy-to-learn examples, which are deemed redundant and removed while pruning. This can lead to poor data coverage at high pruning factors, as only difficult-to-learn examples are retained, causing severe performance drop [68, 77]. CCS, a coverage-driven method, overcomes this limitation by proposing a probability density-based distribution coverage strategy to improve the coverage of the coreset [68]. CCS is adopted here due to its exceptional performance at high pruning factors and its flexibility in incorporating any importance score.
@@
Through the use of these data-pruning techniques, we explore the degree of redundancy in electronic structure data across a wide variety of material systems — a metal, a chemically complex disordered alloy system, and molecules. We show that high degrees of redundancy appear to be a generic feature of electronic structure data: even random pruning can substantially reduce dataset size with minimal loss in predictive accuracy, while state-of-the-art coverage-based methods unveil even more dramatic levels of redundancy. Thus, a remarkably large fraction of the dataset across all the materials studied can be systematically pruned while preserving prediction accuracy and model generalizability. We have further quantified the quality of predictions via calculations of the ground-state energy from predicted electron densities and have found that chemically accurate predictions (i.e., energy errors less than $1$ kcal/mole or $1.6$ mHa/atom) can be made even after pruning $99$% of the data. In particular, our results for complex disordered metallic alloys hold across composition space, thus suggesting the potential to identify a minimal, essential dataset representative of each material class. Additionally, we found that for the materials considered here, ML model training time on the pruned datasets can be reduced by a factor of three or more, thus enabling accelerated hyperparameter tuning and foundation model development.
@@ II Results
Figure 1: Error in ML-predicted electron density for the original dataset, randomly pruned dataset, and 90% and 99% CCS-based pruned dataset.
Figure 2: Error in energy obtained from the ML-predicted electron density for the original dataset, 90% and 99% CCS-based pruned dataset.
@@
In this section, we quantify the redundancy in electronic structure data for three different materials. We present the error in the ML model for various pruning factors to identify the fraction of data that can be pruned without significantly compromising the model’s performance and generalizability, and further quantify this through the accuracy of the energy computed from the ML-predicted electron density.
@@
The dataset consists of snapshots of KS-DFT simulation taken at intermittent time steps, with each snapshot capturing the atomic coordinates, species information and the resulting electron density over the simulation cell. Given a grid point in the simulation domain, the ML model provides a map to the electron density at that point, from the nearby atomic configuration. Details of the KS-DFT simulations and the ML map are provided in Section IV.1.
To explore the redundancy, pruning is performed by removing a subset of the grid points of the original dataset. We explore three approaches for data pruning: random grid-point-wise pruning, GraNd score-based pruning, and CCS-based pruning. In random grid-point-wise pruning, grid points from the original dataset are removed randomly. The importance score (i.e. GraNd score) is computed for each datapoint and used for the GraNd and CCS approaches. Upon training ML models on the pruned datasets, the errors in the predicted electron density on a testing dataset are computed. We report the commonly used normalized root mean squared error (NRMSE) [30] as and error metric, defined in the supplemental materials V.1.
Figure 1 highlights the redundancy in electronic structure data. For all three systems, $90\%$ of the data can be pruned through the CCS-method, without compromising accuracy in the predicted electron density. Even $99\%$ redundancy can be identified by the CCS method: at such a high pruning factor, the errors increase as expected, but the predicted electron densities still result in chemically accurate energies (i.e., errors less than $1$ kcal/mole or $1.6$ mHa/atom) across all system, as shown in Figure 2. Notably, while the errors in the electron density and the corresponding energy errors are correlated, the one-to-one relationship between them can break down when the density errors and hence the energy errors are themselves low. This situation can arise for low pruning percentages across all models, or for very high pruning percentages, for some models. An additional issue to keep in mind is that in general, the CrFeCoNi alloy system is more difficult to predict accurately through ML models, because of the involvement of harder pseudpotentials and semi-core electrons for transition metals in the KS-DFT calculations, and the large amount of compositional variation [28]. Thus, the overall ML predictions are slightly less accurate for this system, as opposed to the other systems studied here. However, our key finding, in this regard is that there is still a high degree of redundancy in the CrFeCoNi electronic structure data, and the elimination of this redundancy does not reduce ML model prediction accuracy further.
@@
Figure 3: (a) $H^{1}$ seminorm and (b) $H^{1}$ norm of the error for the ML-predicted electron density obtained from the original, $90\%$ CCS and $99\%$ CCS based pruned dataset.
@@
To demonstrate that the ML models not only capture the electron density field but also smooth spatial variations in it, we plot the $H^{1}$ seminorm and $H^{1}$ norm of the error in electron density field in Figure 3 (see the supplemental materials V.1 for further details). We see a trend very similar to what is seen in Figure 1. For all the three material systems, the $90\%$ CCS dataset achieves $H^{1}$ seminorm and $H^{1}$ norm values similar to the original dataset, whereas $99\%$ CCS dataset shows a slight increase. The $90\%$ and $99\%$ random pruning dataset has a higher $H^{1}$ seminorm and $H^{1}$ norm than CCS based pruning for all three systems.
In the following, we present the detailed performance of the pruning techniques for various pruning factors using aluminum, CrFeCoNi, and water as representative examples of a metal, a complex alloy system, and molecule, respectively.
II.1 Aluminum
We find that just $10\%$ of the data for Aluminum systems can be selected through the CCS method to match the performance of the original dataset. Moreover, even $1\%$ of the data can be selected to yield a remarkably accurate ML model, that maintains chemical accuracy in the energies.

### v2

@@ Abstract
Accurate prediction of electronic structure underpins advances in chemistry, materials science, and condensed matter physics. In recent years, Machine Learning (ML) has enabled the development of powerful surrogate models that can enable the prediction of the ground state electron density and related properties at a fraction of the computational cost of conventional first principles simulations. Such ML models typically rely on massive datasets generated through expensive Kohn-Sham Density Functional Theory calculations. A key reason for relying on such large datasets is the lack of prior knowledge about which portions of the data are essential, and which are redundant. This study reveals significant redundancies in electronic structure datasets across various material systems, including molecules, simple metals, and chemically complex alloys — challenging the notion that extensive datasets are essential for accurate ML-based electronic structure predictions. We demonstrate that even random pruning can substantially reduce dataset size with minimal loss in predictive accuracy. Furthermore, a state-of-the-art coverage-based pruning strategy that selects data across all learning difficulties, retains chemical accuracy and model generalizability using up to $100$-fold less data, while reducing training time by threefold or greater. By contrast, widely used importance-based pruning methods, which eliminate easy-to-learn data, can catastrophically fail at higher pruning factors due to significant reduction in data coverage. This heretofore unexplored high redundancy in electronic structure data holds the potential to identify a minimal, essential dataset representative of each material class.
@@ I Introduction
The calculation and analysis of the electronic structure of materials underpins much of modern computational materials science. From predicting thermodynamic stability and charge transport to modeling chemical reactivity, the electronic structure serves as a fundamental description of how interacting atoms and electrons give rise to observable properties. Consequently, electronic structure calculations often serve as an important building block of multiscale materials modeling [tadmor2011modeling] with a significant share of contemporary high performance computing resources being devoted to such simulations [hafner2008ab, curtarolo2013high, lejaeghere2016reproducibility].
@@
Kohn-Sham Density Functional Theory (KS-DFT) serves as the primary workhorse of modern electronic structure calculations [kohn1965self, hohenberg1964inhomogeneous, hafner2006toward]. Although formulated in terms of the so-called Kohn-Sham orbitals, the fundamental unknown in KS-DFT is the electron density. Indeed, many ground state material properties, e.g. structural parameters and elastic constants, may be computed from the electron density field, and it also serves as a starting point for ab initio descriptions of excited state phenomena. However, practical calculations using KS-DFT, especially for complex systems such as disordered alloys or materials with extended defects, can be limited by the cubic scaling computational cost of the method with respect to system size. As such, a considerable share of research has been devoted to the development of specialized solution techniques and high performance computing algorithms to overcome this bottleneck [gavini2023roadmap, goedecker1999linear, banerjee2016chebyshev, banerjee2018two, motamarri2014subquadratic, lin2014siesta, dogan2023real, banerjee2021ab, banerjee2016cyclic].
@@
In recent years, Machine Learning (ML) has enabled the development of powerful surrogate models that can enable the prediction of the ground state electron density and related properties at a fraction of the computational cost of conventional first principles simulations [lewis2021learning, jorgensen2022equivariant, zepeda2021deep, chandrasekaran2019solving, fiedler2023predicting, brockherde2017bypassing, del2023deep, schleder2019dft, kulik2022roadmap, pathrudkar2022machine, pathrudkar2024electronic, pathrudkar2024HEA]. These ML models are typically trained on KS-DFT data and use descriptors of the atomic environment to produce the electron density as the output. Recent studies have demonstrated the success of such ML models in predicting electron densities across a wide range of systems, including pure metals, alloys, insulators, organic molecules and nanostructures  [zepeda2021deep, li2025image, chandrasekaran2019solving, pathrudkar2024electronic, pathrudkar2024HEA, pathrudkar2022machine]. Moreover, these models show promising extrapolation capabilities, accurately predicting electron densities for unseen compositions in complex alloys, atomic configurations with defects and dislocations, and for chemical species beyond training [li2025image, pathrudkar2024HEA, pathrudkar2024electronic]. The ML models have also been continuously improved to address several challenges, including the enforcement of physical symmetries, incorporation of uncertainty quantification capabilities, and improvements in training and prediction efficiency [jorgensen2022equivariant, teh2021machine, pathrudkar2024electronic].
@@
The paradigm shift brought on by the development of ML based electronic structure techniques has motivated the creation of extensive datasets [ramakrishnan2014quantum, nandi2023multixc, pinheiro2020machine, liang2019qm] — sometimes comprising millions of structures — under the implicit assumption that large, exhaustive sampling is essential to achieve high predictive accuracy and generalizability [faber2017prediction, smith2018less, musil2019fast]. While this assumption has driven progress in materials informatics [jain2013commentary, saal2013materials, jain2020materials], it also raises questions about the efficiency and necessity of such data collection efforts, particularly given the computational cost and environmental footprint associated with high-throughput quantum mechanical calculations.
@@
In this work, we examine the following fundamental but largely overlooked questions: How much of this electronic structure data is genuinely essential? Furthermore, how can such essential datasets be systematically obtained? To answer these questions, we methodically probe the redundancy in electron density data spanning molecular systems, elemental metals, and chemically complex concentrated alloys, and show that a surprisingly small subset is often sufficient to train ML models to within chemical accuracy. In the context of this work, we define redundancy as the presence of data points whose removal does not significantly impact ML model accuracy, generalizability, or derived physical quantities. Our findings challenge the prevailing view that predictive fidelity and generalizability require vast amounts of data and suggest that it is possible to identify minimal datasets that dramatically reduce computational cost without sacrificing accuracy. Thus our exploration of the redundancy of electronic structure data can potentially lead to a small foundational dataset for each material class, which in turn can be used for benchmarking and training foundation models.
@@
The exploration of redundancy in chemical datasets has received prior attention in the literature [li2023exploiting, li2024md, chen2024beyond]. For instance Li et. al. [li2023exploiting] have demonstrated a high degree of redundancy in datasets of scalar material properties such as the formation energy and band gaps, and have proposed an iterative strategy to reduce the size of training data. However, there has not been a systematic exploration of redundancies in the electronic structure itself, particularly in the context of ML models. Notable prior efforts include the observation that electronic structure data for various simple systems often lie on low-dimensional manifolds, enabling reconstruction via low-rank tensors [blesgen2012approximation, motamarri2016tucker] or a small number of principal component modes [teh2021machine, pathrudkar2022machine]. Somewhat relatedly, recent contributions have also exploited redundancies in the electron density to develop strategic sampling schemes for grid-based representations [feng2025efficient, focassio2023linear], although the effectiveness of such approaches may depend sensitively on the manner in which core states are modeled. Thus, although previous studies have addressed redundancy in various contexts, statistically grounded one-shot pruning methods — free from ad hoc heuristics — have yet to be explored for electronic structure data. Accordingly, determining a statistically justifiable level of data redundancy remains an open problem, which we address here. Thus, the techniques explored here are applicable to any class of material without the need for domain knowledge and possibly even to fields beyond the electron density, e.g. the electronic wavefunctions or the local density of states.
@@
The observation of redundancy in electronic structure data is closely connected to Kohn’s nearsightedness principle [prodan2005nearsightedness, kohn1996density]. Indeed, this result has been used to deduce that the single particle density matrix that appears in electronic structure calculations can be very sparse [benzi2007decay, benzi2013decay], and has led to a large number of electronic structure codes that exploit it [gavini2023roadmap, baer1997sparsity, suryanarayana2017nearsightedness, mohr2015accurate, shimojo2001linear, SURYANARAYANA2018288, nakata2020large, garcia2009linear, goedecker1999linear, skylaris2005introducing]. In the realm of ML-based electronic structure predictions, similar considerations can be used to motivate the prunability of electronic structure data: since local environments largely determine the electronic structure, many configurations differing only in distant parts of the system are expected to have nearly identical electronic descriptors. This implies substantial redundancy in datasets that exhaustively enumerate configurations, especially in large and complex systems. However, there has not been a systematic exploration of this aspect, particularly in the context of state-of-the-art neural network based ML models of electronic structure.
@@
Neural network performance generally improves with data and model size, often following a power law scaling [kaplan2020scaling, hestness2017deep]. However, beyond a certain level, marginal accuracy gains require significantly more data [sorscher2022beyond, zhai2022scaling]. Thus, power law scaling suggests that much of the training data is redundant and can be pruned without notable performance loss [sorscher2022beyond].Commonly used approaches for selecting a smaller, representative subset of a large dataset that preserves ML model performance are referred to as data pruning techniques or coreset selection methods. These techniques are often broadly categorized into the following three major approaches [tan2024data], importance score based methods [paul2021deep, toneva2018empirical, tan2024data, he2024large], coverage driven methods [zheng2023_CCS, sener2018coreset, xia2022moderate] and optimization-based methods [yang2022dataset, mirzasoleiman2020coresets, killamsetty2021grad, killamsetty2021glister, killamsetty2021retrieve].
@@
Most existing pruning methods have been evaluated primarily on classification problems, whereas electronic structure prediction is a regression task. Therefore, identifying the degree of prunability of electronic structure data, as well as the most suitable pruning technique remain unexplored. In the present work, two state-of-the-art methods: GraNd [paul2021deep] and Coverage-Centric Coreset Selection (CCS) [zheng2023_CCS] are adopted. Optimization-based methods are not considered in this work because, despite their strong theoretical foundation and impressive performance, they are computationally quite expensive, and can be challenging to implement on large-scale datasets [yang2024mind]. The GraNd technique is adopted because of its intuitive approach, ease of implementation, and ability to identify important examples early in the training. GraNd, an importance-score based method, assigns lower scores to easy-to-learn examples, which are deemed redundant and removed while pruning. This can lead to poor data coverage at high pruning factors, as only difficult-to-learn examples are retained, causing severe performance drop [zheng2023_CCS, guo2022deepcore]. CCS, a coverage-driven method, overcomes this limitation by proposing a probability density-based coverage strategy to improve the coverage of the coreset [zheng2023_CCS]. CCS is adopted here due to its exceptional performance at high pruning factors and its flexibility in incorporating any importance score.
@@
Through the use of these data-pruning techniques, we explore the degree of redundancy in electronic structure data across a wide variety of material systems — a simple metal, chemically disordered concentrated complex alloy systems, and molecules. We show that high degrees of redundancy appear to be a generic feature of electronic structure data: even random pruning can substantially reduce dataset size with minimal loss in predictive accuracy, while state-of-the-art coverage-based methods unveil even more dramatic levels of redundancy. Thus, a remarkably large fraction of the dataset across all the materials studied can be systematically pruned while preserving prediction accuracy and model generalizability. We have further quantified the quality of predictions via calculations of the ground-state energy from predicted electron densities and have found that chemically accurate predictions (i.e., energy errors less than $1$ kcal/mol or $1.6$ mHa/atom) can be made even after pruning 9999% of the data. In particular, our results for complex disordered metallic alloys hold across composition space, thus suggesting the potential to identify a minimal, essential dataset representative of each material class. Additionally, we found that for the materials considered here, ML model training time on the pruned datasets can be reduced by a factor of three or more, thus enabling accelerated hyperparameter tuning and foundation model development.
@@ II Results
In this section, we quantify the redundancy in electronic structure data for four different materials. We present the error in the ML model for various pruning factors to identify the fraction of data that can be pruned without significantly compromising the model’s performance and generalizability, and further quantify this through the accuracy of the energy computed from the ML-predicted electron density.
The four material systems considered here are pure aluminum, bulk water, and two different alloy systems across composition space — SiGeSn and CrFeCoNi. The first two serve as representative examples of simple metals and molecular systems, respectively, while the latter two illustrate technologically important chemically complex concentrated alloys. Notably our results for the alloy systems (discussed below) are particularly striking since the ML predictions are carried out over the entire composition space of each alloy [pathrudkar2024electronic].
@@
The dataset consists of snapshots of KS-DFT simulation taken at intermittent time steps, with each snapshot capturing the atomic coordinates, species information, and the resulting electron density over the simulation cell. Given a grid point in the simulation domain, the ML model provides a map to the electron density at that point, from the nearby atomic configuration. Details of the KS-DFT simulations and the ML map are provided in Section IV.1.
II.1 Redundancy in diverse material datasets
@@
Figure 1: Error in ML-predicted electron density for the original dataset, 90% and 99% randomly, and CCS-based pruned datasets. Each ML model was trained three times and the mean error is reported.
Figure 2: Error in energy with respect to KS-DFT, as obtained from the ML-predicted electron density for the original dataset, 90% CCS, and 99% CCS-based pruned datasets. The electron density prediction from one of the ML models was postprocessed in each case.
To explore the redundancy, pruning is performed by removing a subset of the grid points of the original dataset. We explore three approaches for data pruning: random grid-point-wise pruning, GraNd score-based pruning, and CCS-based pruning. In random grid-point-wise pruning, grid points from the original dataset are removed randomly. The importance score (i.e. GraNd score) is computed for each datapoint and used for the GraNd and CCS approaches. Upon training ML models on the pruned datasets, the errors in the predicted electron density on a testing dataset are computed. We report the commonly used normalized root mean squared error (NRMSE) [teh2021machine] as an error metric, defined in the supplemental materials (section V.1).
Figure 1 highlights the redundancy in electronic structure data. For all four systems, $90\%$ of the data can be pruned through the CCS-method, without compromising accuracy in the predicted electron density. Even $99\%$ redundancy can be identified by the CCS method: at such a high pruning factor, the errors increase as expected, but the predicted electron densities still result in chemically accurate energies (i.e., errors less than $1$ kcal/mole or $1.6$ mHa/atom) across all system, as shown in Figure 2. The total energy of the system is calculated from the ML-predicted electron density, via postprocessing [pathrudkar2024HEA]. Notably, while the errors in the electron density and the corresponding energy errors are correlated, the one-to-one relationship between them can break down when the density errors and hence the energy errors are themselves low. This situation can arise for low pruning percentages across all models, or for very high pruning percentages, for some models (e.g. those involving the alloy systems considered in this study). An additional issue to keep in mind is that in general, the CrFeCoNi alloy system is more difficult to predict accurately through ML models, because of the involvement of harder pseudopotentials, the inclusion of semi-core electrons for transition metals in the KS-DFT calculations, and the large amount of compositional variation [pathrudkar2024HEA]. Thus, the overall ML predictions are slightly less accurate for this system, as opposed to the other systems studied here. However, our key finding, in this regard is that there is still a high degree of redundancy in the CrFeCoNi electronic structure data, and the elimination of this redundancy does not reduce ML model prediction accuracy further.
Figure 3: (a) $H^{1}$ seminorm and (b) $H^{1}$ norm of the error for the ML-predicted electron density obtained from the original dataset, $90\%$, and $99\%$ randomly and CCS based pruned dataset.
Figure 4: Two-dimensional (2D) slices showing electron density obtained by KS-DFT, ML model trained on the original dataset, 90% CCS, and 99% CCS based pruned dataset for (a) Aluminum at 1500K (b) Water at 600K (c) SiGeSn at 2400K (d) CrFeCoNi for 5000K. The electron density($\rho$) along the solid line is also compared with the KS-DFT and the ML-predicted electron densities show remarkable agreement with the KS-DFT, true for all three ML models across four systems. The unit of electron density is $\text{Bohr}^{-3}$ in atomic units.
@@
To demonstrate that the ML models not only capture the electron density field but also smooth spatial variations in it, we plot the $H^{1}$ seminorm and $H^{1}$ norm of the error in electron density field in Figure 3 (see the supplemental materials, section V.1 for further details). The errors in smoothness exhibit a trend similar to that of the electron density shown in Figure 1. For all four material systems, the $90\%$ CCS-pruned dataset achieves a similar degree of spatial smoothness as the original dataset (i.e., similar $H^{1}$ seminorm and $H^{1}$ norm values), whereas the $99\%$ CCS-pruned dataset shows a slight decrease. The $90\%$ and $99\%$ randomly pruned dataset has higher $H^{1}$ seminorm and $H^{1}$ norm values than CCS based pruning, for all four systems.
We compare the electron density fields obtained from KS-DFT with those predicted by ML models trained on three datasets — the original dataset, the $90\%$, and $99\%$ CCS-based pruned dataset — for all four material systems, as shown in Figure 4. ML models trained on the pruned datasets show excellent agreement with those trained on the original dataset and with KS-DFT calculations across all systems, confirming the high degree of lossless pruning. More details on this comparison are provided in the supplemental materials (section V.7).
II.2 Comparative assessment of pruning strategies